40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 19 18:10:10 2022, Paco, Update, BHD, Part V of BHR upgrade - POX11 path, LO path, and ITMX Oplev DJI_0167.JPGDJI_0169.JPGDJI_0168.JPG
    Reply  Wed Apr 20 08:20:22 2022, Paco, Update, BHD, Part V of BHR upgrade - PR2 weirdness 
       Reply  Wed Apr 20 18:50:03 2022, Koji, Update, BHD, Part V of BHR upgrade - PR2 weirdness Screenshot_2022-04-20_18-46-11.png
    Reply  Thu Apr 21 16:49:01 2022, Paco, Update, BHD, Part V of BHR upgrade - BS Oplev + LO1 offset 
       Reply  Thu Apr 21 18:42:47 2022, Tega, Update, BHD, Part V of BHR upgrade - Align LO and AS beams to BHD BS 
          Reply  Thu Apr 21 20:33:31 2022, Koji, Update, BHD, AS1 UR OSEM problem localized in the chamber Screenshot_from_2022-04-21_20-13-13.png
             Reply  Fri Apr 22 12:03:09 2022, Tega, Update, BHD, AS1 UR OSEM problem localized in the chamber 
                Reply  Fri Apr 22 12:15:08 2022, Anchal, Update, BHD, Cable post installation 
                   Reply  Mon Apr 25 16:57:57 2022, Anchal, Update, BHD, Cable post installation 
Message ID: 16800     Entry time: Thu Apr 21 18:42:47 2022     In reply to: 16797     Reply to this: 16801
Author: Tega 
Type: Update 
Category: BHD 
Subject: Part V of BHR upgrade - Align LO and AS beams to BHD BS 

[Tega, Yuta, Paco]

We tried aligning the LO and AS beams on to the BHD beamsplitter. During the alignment process, we noticed that the damping loop for AS1 was not working. Paco drew our attention to the fact that the UR OSEM signal was alway close to zero, so we checked to ensure that the magnet was still within the OSEM recess and it looks OK. Next we checked the electrical connection at the interface between the copper OSEM cables to the blue in-vacuum flat cable and this too looks alright also. Since the AS1 coil driver was recently modified, it is possible we might find the problem there, so I'll ask Koji about this.


So Koji clarified that the coil driver board and SATAMP boards are different so we should connect this issue to the coil driver board.

ELOG V3.1.3-